Find its inverse
(a) g(x)=(x-1)/(3x+2)
(b) f(x)=4+ square root of (x^2+1)
1
Expert's answer
2020-11-09T17:58:35-0500

"(a)\\ g(x)=\dfrac{(x-1)}{(3x+2)} \implies (3x+2)g(x)=x-1\\
\implies3xg(x)+2g(x)-x+1=0 \implies 2g(x)+1=x-3xg(x) \\
\implies x=\dfrac{1+2g(x)}{1-3g(x)} \\
\\\implies\ g^{-1}(x)=\dfrac{1+2x}{1-3x}; x{=}\mathllap{/\,}1/3\\ \\

(b) \\ f(x)=4+\sqrt{\smash[b]{x^2+1}}; R(f):f\ge5 \implies f(x)-4=\sqrt{x^2+1}\\
\implies (f(x)-4)^2=x^2+1\\ 
\implies (f(x)-4)^2-1=x^2\\
\implies x=\sqrt{(f(x)-4)^2-1}\\
f^{-1}(x)=\sqrt{{(x-4)^2-1}}; x\ge5"